OpenStudy (anonymous):

Please can any one make me understand 'bout HIGH level probability question...? *Conditions are : Step wise Step problem solving : and reasoning

Parth (parthkohli):

Do you mean combinatorics?

OpenStudy (anonymous):

ya similar to it

Parth (parthkohli):

Take an example: Probability of getting 1 head in 5 tosses.

OpenStudy (anonymous):

okay

Parth (parthkohli):

Or wait, let's take this one: Probability of getting at least 1 head in 5 tosses.

OpenStudy (anonymous):

okay its same

Parth (parthkohli):

1) Getting at least 1 head means getting no tails. 2) P(Getting no tails) = 1 - P(Getting all tails)

OpenStudy (anonymous):

kk

Parth (parthkohli):

P(Getting all tails) = 1/2 * 1/2 * 1/2 * 1/2 * 1/2 = 1/32 right?

OpenStudy (anonymous):

k

Parth (parthkohli):

1 - 1/32 = 32/32 - 1/32 = 31/32

OpenStudy (anonymous):

okay

Parth (parthkohli):

Not getting all tails = 31/32 = Getting at least one head.

Parth (parthkohli):

So, the final probability is 31/32. Did you understand?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanx

Parth (parthkohli):

You're welcome!

OpenStudy (anonymous):

np

OpenStudy (lgbasallote):

@Mimi_x3 is who you want lol